Which proportion below can be solved to find h ow many tons are in 5,240,000?
The formula is 1/t = 2000/5240000. I know what the answer is, but where does the 1 come from?

Answer by stanbon(48568)

1/t = 2000/5240000
The number of tons is really 524000/2000
For some reason they have given the problem
in an inverse form.
Instead of t=5240000/2000
they have given you 1/t=2000/5240000.
Cheers,
Stan H.